Attendees: Bruce Glagola, Bruno Fieramosca, Clay White, Nena Moonier, Wendy VanWingeren, Shane Flood, Dean Wyncott, and Patti Pedergnana
- There was an autoinfo email sent by George Doktorcyk regarding the switchgear maintenance during shutdown, which will affect the chilled and process water in all buildings. Bruce asked the group to send a reminder to all of our residents.
- Bruce will be out of the office all day Wednesday and on Friday afternoon. Patti will attend the Shutdown Planning meeting on Wednesday.
- Dean thanked everyone for helping cover his shifts and area while he was out recovering from his shoulder surgery.
- SI completed termination of the emergency egress cables at 16-BM-C. Patti will put the beamline back global online and grant APS enable to the A station.
- Steve Davey provided verbal approval for the Sector 22 mirror tank project. It does not need design review per Davey. Shane will request the approval in writing.
- Sector 2 had a shutter pressure switch replacement today and MOM group requested a CCWP. The group agreed to post a CCWP for any work in the shutter pneumatics boxes.
- 6-ID-C’s radioactive sample experiment ended today but the samples are in the cryostat and need to warm up before coming out of the station. The samples should come out tomorrow before noontime.